a car travelling at 50km/h completes a journey in 6 hours. how long will it take traveling a 75km/h to complete the same journey?
Answer by stanbon(75887)

a car traveling at 50km/h completes a journey in 6 hours.
distance = r*t = 50*6 = 300 km
----------

how long will it take traveling a 75km/h to complete the same journey?
t = distance /rate
---
t = 300/75 km/h = 4 hours
==============================
Cheers,
Stan H.